OpenStudy (jessiegonzales):

MEDAL ANYONE HELP!!! Multiply. (2n+2)(2n-2)

OpenStudy (anonymous):

(a+b)(a-b) = a^2-b^2

OpenStudy (anonymous):

just subsitute for that formula, 2n would be a and 2 would be b

OpenStudy (jessiegonzales):

4N^2+4 @joanip_

OpenStudy (anonymous):

no, just 2n^2-4

OpenStudy (jessiegonzales):

WELL THATS NO ONE OF MY OPTIONS 4N^2-4 4N^2-4N-4 4N^2+2N-4 4N^2+4N-4 @JOANIP_

rishavraj (rishavraj):

its of type \[(a ^{2}-b ^{2}) = (a+b) (a-b)\]

OpenStudy (anonymous):

 It's better to learn how to do it rather than looking at a formula.